HUBBLE SPACE TELESCOPE DAILY REPORT #2420 PERIOD COVERED: 0000Z (UTC) 07/22/99 - 0000Z (UTC) 07/23/99 Daily Status Report as of 204/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.2 Completed STIS/CCD/MA1 7345 (Origin and Physical Conditions in the Local Lyman-Alpha Forest) The Space Telescope Imaging Spectrograph (MA1) was used to make observations of PG1211+143 in order to study the origin and physical conditions in the local Lyman-Alpha forest. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 8444 (WF/PC-2 CYCLE-8 Internal Monitor) The WF/PC-2 was used to make a bias, and an internal flat, observation as part of a Cycle-8 monitoring program to monitor the health of the cameras. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed WF/PC-2 8195 (The Stellar Halo and the Metallicity Distribution Function in the Giant Elliptical NGC 5128) The WF/PC-2 was used to make observations of the giant elliptical galaxy NGC-5128 in order to study its stellar halo and the metallicity distribution function. Because this galaxy is 2 magnitudes closer than any other giant elliptical galaxy, it gives us a unique opportunity to study the enrichment history at a level of detail that is impossible for any other such galaxy. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.8 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.9 Completed STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ0200-3 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 8037 (STIS Coronographic Imaging of Circumstellar Discs) The Space Telescope Imaging Spectrograph (CCD) was used to make coronographic observations of the circumstellar protoplanetary disk around the star Fomalaut (a.k.a., Alpha PsA). The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 7374 (Probing Low Density Cluster Environments at Moderate-to-High Redshifts) The WF/PC-2 was used to make observations of the galactic cluster Gal-Clus-0723. Recent HST WF/PC-2 images of rich clusters at moderate-to-high redshifts have revealed that morphological information is a crucial ingredient in understanding the physical processes responsible for the evolution of the galaxy population in clusters. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Two Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Operations Notes: Upon conclusion of GSFC Internal Simulation #8 (EVA-3), the operations team returned to the STOCC. No problems were encountered while operating out of the SAMS back-up facility. Using ROP SR-1, the SSR EDAC error counter was cleared three times. HSTAR 7201 documents limit violations for the SA wing blanket temperature 2 at 203/180959Z (5 minutes) and at 203/1994823Z (3 minutes). Each low reading (-102.1 and 104.0 respectively) was at the end of the orbit night period. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH